# Leptogenesis and Neutrino Masses Via Pseudo-Dirac Gauginos

### v1

@@ V Signals
The generation of the BAU imposes strong requirements on the parameter space of this model. We have focused on the bino and the wino with masses $\mathcal{O}({\rm TeV})$. In order to have a long-lived bino that decays out of equilibrium, we need the bino to be either the lightest neutralino or at least $M_{\tilde{B}}<M_{\widetilde{W}}+M_{Z}$ and a high messenger scale, $\Lambda_{M}\gtrsim 10^{7}$ TeV. In Gehrlein and Ipek (2021), the LHC phenomenology of a long-lived bino was studied in a $U(1)_{R-L}$-symmetric MSSM model like this one, for messenger scales of up to $10^{11}$ TeV with signals inside the ATLAS detector or at proposed detectors such as MATHUSLA Curtin et al. (2019); Alpigiani et al. (2022) and CODEX-b Gligorov et al. (2018); Aielli et al. (2022). However, those results were obtained for a scenario in which binos were produced in squark decays. In this leptogenesis scenario, out-of-equilibrium conditions require heavy sfermions with masses $M_{\rm sf}\gtrsim 25$ TeV, which are inaccessible at the LHC.
@@
    $M_{\widetilde{W}}+M_{Z}>M_{\tilde{B}}>M_{\widetilde{W},\chi^{\pm}}$: In this case, the bino, wino, and charginos will all decay only via their mixing with neutrinos and charged leptons, respectively. For prompt wino/chargino decays, an ATLAS search for trilepton resonances with $139~{}{\rm fb}^{-1}$ data excludes wino/chargino with masses between 100 GeV and 1.1 TeV Aad et al. (2021). Here, in contrast to Ayber and Ipek (2023), the bino and wino can both be long-lived due to the large messenger scale. Hence, these constraints might not apply.
@@ VI Conclusions and Outlook
We studied a $U(1)_{R-L}$-symmetric supersymmetric model that can generate both the light neutrino masses and the baryon asymmetry of the universe. It has been shown that the pseudo-Dirac bino and wino in this model can result in a hybrid type I + III inverse seesaw mechanism that explains the neutrino masses Ayber and Ipek (2023). In this work, we separated the two mechanisms: the wino is responsible for the neutrino mass generation while the pseudo-Dirac bino generates the BAU through its out-of-equilibrium decay in the early universe. The leptogenesis scenario we studied requires a bino of mass a few TeV with a small Majorana mass of $1010^{-3}~{}{\rm eV}-$ 10 eV. The smallness of the Majorana gaugino masses can be related to how the $U(1)_{R-L}$ is broken and how this breaking is mediated to the supersymmetric sector.
@@
Pseudo-Dirac fermions can oscillate between their particle and antiparticle states with an oscillation frequency $\omega\sim 2m$ where $m$ is the Majorana mass. If the particle and antiparticle states can decay into the same final state, these oscillations can enhance the $CP$ violation in the system. In this work, we used decays facilitated by bino-neutrino mixing terms in the neutrino mass Lagrangian given in eq. 7. This is a leptogenesis scenario and the bino needs to decay before EW sphalerons turn off at $T_{\rm sph}\simeq 130$ GeV. Together with the out-of-equilibrium condition, successful leptogenesis requires a messenger scale $\Lambda_{M}\sim O(10^{7}~{}{\rm TeV})$ for a bino of mass a few TeV.
@@
The oscillations in the early universe are affected by both the expansion of the universe and the interactions of the particles with the rest of the plasma. Unlike the most generic leptogenesis scenarios where the right-handed neutrinos interact with the SM only through their mixing with the light neutrinos, bino has much more effective interactions as it couples to a sfermion and a fermion. These interactions reduce the resulting baryon asymmetry in two ways. They keep the bino in thermal equilibrium, washing away the asymmetry and delay oscillations resulting in smaller $CP$ violation. Even with these hindrances, a large baryon asymmetry, $\Delta_{B}\sim 10^{-6}$, can be produced for $M_{\rm sf}\gtrsim 25$ TeV and $m\sim 10^{-3}-10$ eV.
@@
Generating the observed BAU induces strong requirements on the mass spectrum of the model, pushing it into a decoupled scenario with very heavy sfermions, $M_{\rm sf}\gtrsim 25$ TeV. Still, the wino can have interesting signatures at the LHC. If the wino is the lightest accessible SUSY particle, it decays to final states with leptons, quarks and missing energy, through its mixing with the SM neutrinos. Since the creation of the BAU requires a high messenger scale, these decays could result in displaced vertices. Furthermore, the branching fractions to different lepton flavors point towards the neutrino mass generation mechanism. These signatures are interesting targets for long-lived particle searches for both the current LHC detectors Aad et al. (2024a, b); Hayrapetyan et al. (2024, 2025) and for proposed detectors like MATHUSLA Curtin et al. (2019); Alpigiani et al. (2022) and CODEX-b Aielli et al. (2022). We leave the collider studies to future work.
@@ Acknowledgements
This work was supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada. SI is grateful for the hospitality of Perimeter Institute where part of this work was carried out. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. This work was supported by a grant from the Simons Foundation (1034867, Dittrich).

### v2

@@ V Signals
The generation of the BAU imposes strong requirements on the parameter space of this model. We have focused on the bino and wino with masses $\mathcal{O}({\rm TeV})$. In order to have a long-lived bino that decays out of equilibrium, we need the bino to be either the lightest neutralino or at least $M_{\tilde{B}}<M_{\widetilde{W}}+M_{Z}$ and a high messenger scale, $\Lambda_{M}\gtrsim 10^{7}$ TeV. In Gehrlein:2021hsk, the LHC phenomenology of a long-lived bino was studied in a $U(1)_{R-L}$-symmetric MSSM model like this one, for messenger scales of up to $10^{11}$ TeV with signals inside the ATLAS detector or at proposed detectors such as MATHUSLA Curtin:2018mvb; MATHUSLA:2022sze and CODEX-b Gligorov:2017nwh; Aielli:2022awh. However, those results were obtained for a scenario in which binos were produced in squark decays. In this leptogenesis scenario, out-of-equilibrium conditions require heavy sfermions with masses $M_{\rm sf}\gtrsim 50$ TeV, which are inaccessible at the LHC.
@@
    $M_{\widetilde{W}}+M_{Z}>M_{\tilde{B}}>M_{\widetilde{W},\chi^{\pm}}$: In this case, the bino, wino, and charginos will all decay only via their mixing with neutrinos and charged leptons, respectively. For prompt wino/chargino decays, an ATLAS search for trilepton resonances with $139~{\rm fb}^{-1}$ data excludes wino/chargino with masses between 100 GeV and 1.1 TeV ATLAS:2020uer. Here, in contrast to Ayber:2023diw, the bino and wino can both be long-lived due to the large messenger scale. Hence, these constraints might not apply.
@@ VI Conclusions and Outlook
We studied a $U(1)_{R-L}$-symmetric supersymmetric model that can generate both the light neutrino masses and the baryon asymmetry of the universe. It has been shown that the pseudo-Dirac bino and wino in this model can result in a hybrid type I + III inverse seesaw mechanism that explains the neutrino masses Ayber:2023diw. In this work, we separated the two mechanisms: the wino is responsible for the neutrino mass generation, whilst the pseudo-Dirac bino generates the BAU through its out-of-equilibrium decay in the early universe. The leptogenesis scenario we studied requires a bino of mass a few TeV with a small Majorana mass of $10^{-4}{\rm~eV}-10$ eV. The smallness of the Majorana gaugino masses can be related to how the $U(1)_{R-L}$ is broken and how this breaking is mediated to the supersymmetric sector.
@@
Pseudo-Dirac fermions can oscillate between their particle and antiparticle states with an oscillation frequency $\omega\sim 2m$ where $m$ is the Majorana mass. If the particle and antiparticle states can decay into the same final state, these oscillations can enhance the $CP$ violation in the system. In this work, we used decays facilitated by bino-neutrino mixing terms in the neutrino mass Lagrangian given in eq. 7. This is a leptogenesis scenario, and the bino needs to decay before EW sphalerons turn off at $T_{\rm sph}\simeq 130$ GeV. Together with the out-of-equilibrium condition, successful leptogenesis requires a messenger scale $\Lambda_{M}\sim\mathcal{O}(10^{8}~{\rm TeV})$ for a bino of mass a few TeV.
@@
The oscillations in the early universe are affected by both the expansion of the universe and interactions between particles and the rest of the plasma. Unlike the most generic leptogenesis scenarios, where the right-handed neutrinos interact with the SM only through their mixing with the light neutrinos, bino has much more effective interactions as it couples to a sfermion and a fermion. These interactions reduce the resulting baryon asymmetry in two ways. They keep the bino in thermal equilibrium, washing away the asymmetry and delaying oscillations, resulting in smaller $CP$ violation. Even with these hindrances, a large baryon asymmetry, $\Delta_{B}\sim 10^{-7}$, can be produced for $M_{\rm sf}\gtrsim 100$ TeV and $m\sim 10^{-4}$ eV.
@@
Generating the observed BAU imposes stringent constraints on the mass spectrum of the model, pushing it toward a decoupled scenario with very heavy sfermions, $M_{\rm sf}\gtrsim 50$ TeV. Still, the wino can have interesting signatures at the LHC. If the wino is the lightest accessible SUSY particle, it decays to final states with leptons, quarks, and missing energy through its mixing with the SM neutrinos. Since the creation of the BAU requires a high messenger scale, these decays could result in displaced vertices. Furthermore, the branching fractions to different lepton flavors point towards the neutrino mass generation mechanism. These signatures are interesting targets for long-lived particle searches for both the current LHC detectors ATLAS:2024qoo; ATLAS:2024ocv; CMS:2024qxz; CMS:2024xzb and for proposed detectors like MATHUSLA Curtin:2018mvb; MATHUSLA:2022sze and CODEX-b Aielli:2022awh. We leave the collider studies to future work.
@@ Acknowledgements
This work was supported by the Natural Sciences and Engineering Research Council (NSERC) of Canada and the Arthur B. McDonald Canadian Astroparticle Physics Research Institute. SI is grateful for the hospitality of Perimeter Institute, where part of this work was carried out. Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. This work was supported by a grant from the Simons Foundation (1034867, Dittrich).